# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v4

@@ 6.1.3. Benefits and Limitations
[…]
The illustrations of the side and top view of the camera’s viewable space (in pyramid shape), where points A represents the camera position, and points B, C, D, and E represent the top-left, top-right, bottom-right, bottom-left of the frame at d meters in front of the camera. The left figure shows the camera’s viewable area intersecting with a polygon representing a road intersection. The right figure shows the camera’s viewable area, which does not intersect with any polygon representing a road intersection.
[…]
@@ 6.2.1. High-Level Concepts
Spatialyze only needs to track objects of interest as stated in the workflow. The Object Type Pruner prunes out irrelevant objects from being tracked. We analyze users’ filter predicates to look for types of objects that are necessary for computing the workflow outputs. Using the predicates in § 3 as an example, the Object Type Pruner partially executes the line 3, where users only look for cars and trucks in the workflow outputs. The trajectories of other objects (e.g., humans) will not appear in the final workflow outputs anyway, so tracking them is unnecessary. Therefore, we only need to input objects detected as cars or trucks into the object tracker to reduce its workload.
@@ 6.3.1. High-Level Concepts
0pt leverages Camera to estimate an object’s 3D locations from its 2D bounding box. It is a substitution for Monodepth2 to estimate objects’ 3D locations. 0pt uses basic geometry calculations and requires 2 assumptions to hold. First, the object of interest must be on the ground. We can then assume that the lower horizontal border of the object’s bounding box is where the object touches the ground. Second, we can identify the equation of the plane that the object of interest is on. In Spatialyze, the plane $z=0$ is the ground, although our algorithm can be modified to be used with any plane. The middle point of the lower border of each object’s 2D bounding box represents the point that it touches the ground (red $\star$ in Fig. 3).
@@ 6.3.2. Algorithm
0pt performs 2 steps to estimate the 3D location of an object from its 2D bounding box. First, based on the second assumption, 0pt finds an equation that represents a Vector of Possible 3D Locations of an object. Second, based on the first and second assumptions, 0pt finds an Exact 3D Location from the vector that intersects with the ground.
@@ 6.3.3. Benefits and Limitations
0pt relies on the assumptions mentioned, which do not always hold, especially, the first one. Nevertheless, Spatialyze can analyze users’ filter predicates and tell whether the types of objects of interest can be assumed to touch the ground, applying the optimization only then. For example, “car,” “bicycle,” or “human” touch the ground, while “traffic light” does not. In addition, 0pt is 192$\times$ faster compared than Monodepth2 on average. If one of the estimated 3D locations of an object in a video frame ends up being behind the camera, which means that the object does not touch the ground, Spatialyze can fall back to using Monodepth2 for that frame. Otherwise, 0pt provides 99.48% runtime reduction on average for object 3D location estimation and 52% for the video processing.
[…]
A 3D illustration of geometry-based 3D location estimation, showing relationships between the camera, the 2D car in the image frame, and the actual car location
@@ 6.4.2. Sampling Algorithm
§ 6.4.2 shows the pseudocode of each sampleEvent. exitsLane implements (i) when the car exits the lane. Given the car’s location at the current frame (carLoc) and the lane that contains the car, we assume that the car drives in the lane straight along the lane’s direction. In line 6.4.2, we model the car’s movement as a motion tuple (carLoc, lane.direction). We compute the intersection between the car’s motion tuple and the lane as the location where the car exits the lane. Given the car’s speed v, we know the time it reaches the exit location starting from the current frame in line 6.4.2. We sample the frame before the car reaches the exit location. exitsCamera implements (ii) when the car exits the camera’s view. Knowing the car’s current location, moving direction, and speed, carMoves computes its location at any given frame in line 6.4.2. In line 6.4.2, we use the viewable area of the Camera at each frame derived from § 6.1 to find the first frame when the car’s location is not in the camera’s view. This is the frame when the car already exits the camera’s view, so we sample its preceding frame. newCar implements (iii) when a new car enters the Camera’s view. In line 6.4.2, we sample the next frame that has more cars detected than the current one. The number of cars is a good heuristic to predict that a new car is entering the frame.
@@
{listing}
[!ht]{minted}[linenos, mathescape, escapeinside=!!, fontsize=]python def exitsLane(currentFrame, carLoc, lane, v, cameras): exitLaneLoc = intersection((carLoc, lane.direction), lane)!! exitLaneTime = cameras[currentFrame].time + distance(exitLaneLoc, carLoc)/v!! return lastFrameBefore(exitLaneTime) !! def exitsCamera(currentFrame, carLoc, lane, v, cameras): nextFrame = currentFrame + 1 while cameras[nextFrame].view.contains(carMoves(carLoc, lane.direction, v, !! cameras[nextFrame].time - cameras[currentFrame].time)): nextFrame += 1 !! return lastFrameBefore(nextFrame) def newCar(currentFrame): nextFrame = currentFrame + 1 while numCars(nextFrame) ¡= numCars(currentFrame): nextFrame += 1 !! return nextFrame Computing the events that might trigger a frame sample
@@
First, we compute newCar(1) when new cars enter $Camera_{B}$’s view. In the rest of the videos, only one car is in the frame. We then compute exitsLane(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when $Car_{A}$ exits $Lane_{A}$. In line 6.4.2, $Car_{A}$’s motion tuple intersects the polygon of $Lane_{A}$ at location (66.3, 72.7, 0) at 3:51:49.96 PM in line 6.4.2. The last frame before this time is frame 22, depicted as exitsLane in Fig. 4. Finally, we compute exitsCamera(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when the $Car_{A}$ exits the $Camera_{B}$’s view. The $Car_{A}$ exits the camera view at frame 35 computed in line 6.4.2. Thus, we sample the frame 34, depicted as exitsCamera in Fig. 4.
@@ 6.4.3. Benefits and Limitations
[…]
The figure above is an illustration of exitLane event and exitCamera event. The exitLane event happens earlier. In the figure below, the image shows the first frame that the sampling algorithm starts running on, the frames that the algorithm skips, and the frame that the algorithm skips to (when the car in the first frame exits the lane that it was on). Runtime and Accuracy Evaluation of the Exit Frame Sampler. The left scatter plot shows the object tracking time reduces when the Skip Distance is higher. The right scatter plot shows that the F1 Score (accuracy) of the object tracking algorithm decreases when the Skip Distance is higher. The plot shows our observation that there are more instances of small skipping as a result of the sampling algorithm than instances of large skipping.
[…]
@@ 7. Evaluation
Queries. We implemented Spatialyze’s workflows using four realistic road scenarios, inspired by queries from Scenic (Kim et al., 2021b), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021). These workflows test Spatialyze against real-life situations. Each workflow starts by loading RoadNetworks and Videos with their Cameras from each query’s corresponding datasets, similar to lines 3-3 in § 3. Then, we filter the videos with one of the queries described in Table 1, similar to lines 3-3. All queries look for objects closer than 50 meters. Finally, we save the filtered videos into video snippet files, similar to line 3. Table 1 shows the queries in natural language and S-Flow’s filter predicates.
@@ 7.1.1. EVA
As shown, Spatialyze performs 2-7.3$\times$ faster than EVA on Q5-7, since Spatialyze does not need to run Monodepth2 on each frame, due to the Object Type Pruner and 0pt. In addition, Spatialyze shaves even more time by employing the Road Visibility Pruner to avoid costly YOLO on frames where they are unnecessary. For Q8, our processing time is comparable to EVA. This is because Spatialyze’s Movable Objects Query Engine finds every possible triplet of cars; thus it performs 2 self-joins internally, while EVA only returns video frames with 3 cars or more. We also compare the lines of code (LoC) counts between Spatialyze (14-30) and EVA (58-60), showing that Spatialyze makes query implementation easier and requires less LoC for the same results.
@@ 7.1.5. SkyQuery
[…]
Runtime comparison of Spatialyze against OTIF, NuScenes Devkit, EVA, VIVA (both their data and NuScenes), and SkyQuery (with their data). Spatialyze is faster in all cases except when running Q8 compared to EVA
The bar chart above shows the runtime breakdown of each streaming operator, comparing between SB, S1, S2, S3, S4, S5, S6, and for Q1, Q2, Q3, Q4. The chart below shows the AssA against SB of S1, S2, S4, S5, S6, and for Q1, Q2, Q3, Q4.
[…]
@@ 7.2. Ablation Study
To evaluate each of our optimization techniques, we manually construct video processing plans for the following 7 experiment setups. (SB) Baseline is a Baseline plan without any optimization. It includes the 4 processing operators discussed in § 5.2.2. (S1) is the Baseline plan with the Road Visibility Pruner. (S2) is the Baseline plan with the Object Type Pruner. (S3) is the Baseline plan with the Monocular Depth Estimator replaced with the 0pt. (S4) is the Baseline plan with the Exit Frame Sampler. (S5) enables all optimizations except the Exit Frame Sampler. (S6) enables all optimizations. We evaluated Spatialyze using Q1-4 because Spatialyze can apply all or most of the optimization techniques for the execution of such queries; therefore, we can compare the effects of each optimization technique.
@@ 7.2.1. Execution time
With all optimization techniques, we achieve 2.5-5.3$\times$ speed up over the baseline plan, as shown in Fig. 5. The Road Visibility Pruner affects the entire plan. It prunes out 3.8% of the frames that do not contain “lane” in Q3 and 21.5% of the frames that do not contain “intersection” in Q1-2, with 3.2-20.0% runtime reduction. The Object Type Pruner only affects StrongSORT. It prunes out 36.5% of objects that are not cars in Q2-4 or trucks and 86.3% that are not pedestrians in Q1, with 7.0-26.3% runtime reduction. The 0pt makes the runtime portion of 3D location estimation insignificant (from 48% to 0.55%). Q1 does not execute the Exit Frame Sampler as it only works with cars or trucks. The Exit Frame Sampler has an overhead runtime of executing § 6.4.2; however, it reduces the number of frames into StrongSORT (13.1%), with 0.8% runtime reduction. We also evaluate the execution time with respect to video length by varying the video length input into the workflow. We found that the execution time of the workflow is linearly proportional to video length and Spatialyze can process videos larger than the available memory.
@@ 7.2.2. Accuracy
In Fig. 5, the Road Visibility Pruner (S1) causes a small AssA drop of 0.4% for Q3-4 and 4.7% for Q1-2. The Object Type Pruner (S2) also causes a small AssA drop of 2.5% from Q2-4 and 5.3% for Q1. We do not include (S3) into the results since the object tracker only concerns the 2D bounding boxes of objects; 0pt does not affect the AssA of the object tracker. With all 3 optimization techniques (S5), the video processor still produces accurate object tracks of 93.4% on average compared to (SB), while speeding up 2.5-5.3$\times$ of its runtime. With all optimization techniques (S6), the video processor’s accuracy for object tracking is 84.5% on average, while speeding up 1.65-4.32%, additionally.

### v5

@@ 6.1.3. Benefits and Limitations
Figure 2. (a) A 3D viewable space of a camera in a pyramid shape and its projected area of the pyramid onto the $z=0$ plane. (b) In 1, the video processor executes expensive ML models only on video frames with a visible intersection.
@@ 6.2.1. High-Level Concepts
Spatialyze only needs to track objects of interest as stated in the workflow. The Object Type Pruner prunes out irrelevant objects from being tracked. We analyze users’ filter predicates to look for types of objects that are necessary for computing the workflow outputs. Using the predicates in 1 as an example, the Object Type Pruner partially executes the line 7, where users only look for cars and trucks in the workflow outputs. The trajectories of other objects (e.g., humans) will not appear in the final workflow outputs anyway, so tracking them is unnecessary. Therefore, we only need to input objects detected as cars or trucks into the object tracker to reduce its workload.
@@ 6.3.1. High-Level Concepts
Geometry-Based 3D Location Estimator leverages Camera to estimate an object’s 3D locations from its 2D bounding box. It is a substitution for Monodepth2 to estimate objects’ 3D locations. Geometry-Based 3D Location Estimator uses basic geometry calculations and requires 2 assumptions to hold. First, the object of interest must be on the ground. We can then assume that the lower horizontal border of the object’s bounding box is where the object touches the ground. Second, we can identify the equation of the plane that the object of interest is on. In Spatialyze, the plane $z=0$ is the ground, although our algorithm can be modified to be used with any plane. The middle point of the lower border of each object’s 2D bounding box represents the point that it touches the ground (red $\star$ in Fig. 3).
@@ 6.3.2. Algorithm
Geometry-Based 3D Location Estimator performs 2 steps to estimate the 3D location of an object from its 2D bounding box. First, based on the second assumption, Geometry-Based 3D Location Estimator finds an equation that represents a Vector of Possible 3D Locations of an object. Second, based on the first and second assumptions, Geometry-Based 3D Location Estimator finds an Exact 3D Location from the vector that intersects with the ground.
@@ 6.3.3. Benefits and Limitations
Geometry-Based 3D Location Estimator relies on the assumptions mentioned, which do not always hold, especially, the first one. Nevertheless, Spatialyze can analyze users’ filter predicates and tell whether the types of objects of interest can be assumed to touch the ground, applying the optimization only then. For example, “car,” “bicycle,” or “human” touch the ground, while “traffic light” does not. In addition, Geometry-Based 3D Location Estimator is 192$\times$ faster compared than Monodepth2 on average. If one of the estimated 3D locations of an object in a video frame ends up being behind the camera, which means that the object does not touch the ground, Spatialyze can fall back to using Monodepth2 for that frame. Otherwise, Geometry-Based 3D Location Estimator provides 99.48% runtime reduction on average for object 3D location estimation and 52% for the video processing.
@@ 6.4.2. Sampling Algorithm
3 shows the pseudocode of each sampleEvent. exitsLane implements (i) when the car exits the lane. Given the car’s location at the current frame (carLoc) and the lane that contains the car, we assume that the car drives in the lane straight along the lane’s direction. In line 2, we model the car’s movement as a motion tuple (carLoc, lane.direction). We compute the intersection between the car’s motion tuple and the lane as the location where the car exits the lane. Given the car’s speed v, we know the time it reaches the exit location starting from the current frame in line 3. We sample the frame before the car reaches the exit location. exitsCamera implements (ii) when the car exits the camera’s view. Knowing the car’s current location, moving direction, and speed, carMoves computes its location at any given frame in line 8. In line 7, we use the viewable area of the Camera at each frame derived from § 6.1 to find the first frame when the car’s location is not in the camera’s view. This is the frame when the car already exits the camera’s view, so we sample its preceding frame. newCar implements (iii) when a new car enters the Camera’s view. In line 12, we sample the next frame that has more cars detected than the current one. The number of cars is a good heuristic to predict that a new car is entering the frame.
@@
Listing 3: Computing the events that might trigger a frame sample
1def exitsLane(currentFrame, carLoc, lane, v, cameras):
2 exitLaneLoc = intersection((carLoc, lane.direction), lane)
3 exitLaneTime = cameras[currentFrame].time + distance(exitLaneLoc, carLoc)/v
4 return lastFrameBefore(exitLaneTime)
5def exitsCamera(currentFrame, carLoc, lane, v, cameras):
6 nextFrame = currentFrame + 1
7 while cameras[nextFrame].view.contains(carMoves(carLoc, lane.direction, v,
8 cameras[nextFrame].time - cameras[currentFrame].time)): nextFrame += 1
9 return lastFrameBefore(nextFrame)
10def newCar(currentFrame):
11 nextFrame = currentFrame + 1
12 while numCars(nextFrame) <= numCars(currentFrame): nextFrame += 1
13 return nextFrame
@@
First, we compute newCar(1) when new cars enter $Camera_{B}$’s view. In the rest of the videos, only one car is in the frame. We then compute exitsLane(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when $Car_{A}$ exits $Lane_{A}$. In line 2, $Car_{A}$’s motion tuple intersects the polygon of $Lane_{A}$ at location (66.3, 72.7, 0) at 3:51:49.96 PM in line 3. The last frame before this time is frame 22, depicted as exitsLane in Fig. 4. Finally, we compute exitsCamera(1, (70.92,74.7,0), $Lane_{A}$, 25 mph, $cameras$) when the $Car_{A}$ exits the $Camera_{B}$’s view. The $Car_{A}$ exits the camera view at frame 35 computed in line 8. Thus, we sample the frame 34, depicted as exitsCamera in Fig. 4.
@@ 6.4.3. Benefits and Limitations
Figure 4. (a) Sample Events exitsLane (i) and exitsCamera (ii). (b) Example Results of our Sampling Algorithm. (c) F1 score and runtime reduction for each skip distance when the Exit Frame Sampler can skip at least 1 frame.
@@ 7. Evaluation
Queries. We implemented Spatialyze’s workflows using four realistic road scenarios, inspired by queries from Scenic (Kim et al., 2021b), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021). These workflows test Spatialyze against real-life situations. Each workflow starts by loading RoadNetworks and Videos with their Cameras from each query’s corresponding datasets, similar to lines 1-4 in 1. Then, we filter the videos with one of the queries described in Table 1, similar to lines 5-9. All queries look for objects closer than 50 meters. Finally, we save the filtered videos into video snippet files, similar to line 10. Table 1 shows the queries in natural language and S-Flow’s filter predicates.
@@ 7.1.1. EVA
As shown, Spatialyze performs 2-7.3$\times$ faster than EVA on Q5-7, since Spatialyze does not need to run Monodepth2 on each frame, due to the Object Type Pruner and Geometry-Based 3D Location Estimator. In addition, Spatialyze shaves even more time by employing the Road Visibility Pruner to avoid costly YOLO on frames where they are unnecessary. For Q8, our processing time is comparable to EVA. This is because Spatialyze’s Movable Objects Query Engine finds every possible triplet of cars; thus it performs 2 self-joins internally, while EVA only returns video frames with 3 cars or more. We also compare the lines of code (LoC) counts between Spatialyze (14-30) and EVA (58-60), showing that Spatialyze makes query implementation easier and requires less LoC for the same results.
@@ 7.1.5. SkyQuery
Figure 5. (a) Video frames processed per second for each system. (b) Average video processing time for each video of 20 seconds, comparing each optimization technique. (c) AssA against (SB) vs. other experiment setups.
@@ 7.2. Ablation Study
To evaluate each of our optimization techniques, we manually construct video processing plans for the following 7 experiment setups. (SB) Baseline is a Baseline plan without any optimization. It includes the 4 processing operators discussed in § 5.2.2. (S1) is the Baseline plan with the Road Visibility Pruner. (S2) is the Baseline plan with the Object Type Pruner. (S3) is the Baseline plan with the Monocular Depth Estimator replaced with the Geometry-Based 3D Location Estimator. (S4) is the Baseline plan with the Exit Frame Sampler. (S5) enables all optimizations except the Exit Frame Sampler. (S6) enables all optimizations. We evaluated Spatialyze using Q1-4 because Spatialyze can apply all or most of the optimization techniques for the execution of such queries; therefore, we can compare the effects of each optimization technique.
@@ 7.2.1. Execution time
With all optimization techniques, we achieve 2.5-5.3$\times$ speed up over the baseline plan, as shown in Fig. 5. The Road Visibility Pruner affects the entire plan. It prunes out 3.8% of the frames that do not contain “lane” in Q3 and 21.5% of the frames that do not contain “intersection” in Q1-2, with 3.2-20.0% runtime reduction. The Object Type Pruner only affects StrongSORT. It prunes out 36.5% of objects that are not cars in Q2-4 or trucks and 86.3% that are not pedestrians in Q1, with 7.0-26.3% runtime reduction. The Geometry-Based 3D Location Estimator makes the runtime portion of 3D location estimation insignificant (from 48% to 0.55%). Q1 does not execute the Exit Frame Sampler as it only works with cars or trucks. The Exit Frame Sampler has an overhead runtime of executing § 6.4.2; however, it reduces the number of frames into StrongSORT (13.1%), with 0.8% runtime reduction. We also evaluate the execution time with respect to video length by varying the video length input into the workflow. We found that the execution time of the workflow is linearly proportional to video length and Spatialyze can process videos larger than the available memory.
@@ 7.2.2. Accuracy
In Fig. 5, the Road Visibility Pruner (S1) causes a small AssA drop of 0.4% for Q3-4 and 4.7% for Q1-2. The Object Type Pruner (S2) also causes a small AssA drop of 2.5% from Q2-4 and 5.3% for Q1. We do not include (S3) into the results since the object tracker only concerns the 2D bounding boxes of objects; Geometry-Based 3D Location Estimator does not affect the AssA of the object tracker. With all 3 optimization techniques (S5), the video processor still produces accurate object tracks of 93.4% on average compared to (SB), while speeding up 2.5-5.3$\times$ of its runtime. With all optimization techniques (S6), the video processor’s accuracy for object tracking is 84.5% on average, while speeding up 1.65-4.32%, additionally.